# Dynamical Structure Factor from Weak Measurements

### v1

@@ Abstract
Much of our knowledge of quantum systems is encapsulated in the expectation value of Hermitian operators, experimentally obtained by averaging projective measurements. However, dynamical properties are often described by products of operators evaluated at different times; such observables cannot be measured by individual projective measurements, which occur at a single time. For example, the dynamical structure factor describes the propagation of density excitations, such as phonons, and is derived from the spatial density operator evaluated at different times. Conventionally, this is measured by first exciting the system at a specific wavevector and frequency, then measuring the response. Here, we describe an alternative approach using a pair of time-separated weak measurements, and analytically show that their cross-correlation function directly recovers the dynamical structure factor. We provide numerical confirmation of this technique with a matrix product states simulation of the one-dimensional Bose-Hubbard model, weakly measured by phase contrast imaging. We explore the limits of the method and demonstrate its applicability to real experiments with limited imaging resolution.
@@
The properties of many body quantum systems are often encoded in response functions that quantify the reaction of the system to weak external perturbations. The dynamical structure factor (DSF) $S(q,\omega)$, describing a system’s response to density perturbations with wavenumber $q$ and frequency $\omega$, derives from the expectation value $\langle\hat{n}(0,0)\hat{n}(x,t)\rangle$ of the spatial density operator $\hat{n}(x,t)$ at different times and positions Ashcroft and Mermin (1976). As their name suggests, such quantities are experimentally obtained by observing the system’s response to a suitable perturbation: for example neutron scattering gives access to $S(q,\omega)$ in materials Sturm (1993). Analogously for ultracold atoms, $S(q,\omega)$ can be obtained by Bragg scattering far-detuned laser light off atomic ensembles Steinhauer et al. (2002); Pino et al. (2011); Landig et al. (2015); Biss et al. (2022); the spectral function, a related response function correlating fields rather than densities, can be measured using related techniques Brown et al. (2020); Bohrdt et al. (2020, 2021). By contrast, we focus on measuring density-density correlations $\langle\hat{n}(0,0)\hat{n}(x,t)\rangle$, and thereby $S(q,\omega)$, by weak quantum measurements alone. This connects to previous work Hung et al. (2011), that showed the static structure factor $S(q)$ can be obtained from the same-time density-density correlation function $\langle\hat{n}(0)\hat{n}(x)\rangle$ computed from simple projective measurements of density. Because projective measurements collapse the wavefunction at a well-defined time, they prevent access to the two-time correlations we require.
Figure 1: System and measurement schema. (a) Homodyne detection implemented via PCI. Probe light is first phase shifted by the atomic ensemble (light scattered by more than the acceptance angle $\theta$ is not detected); the un-scattered light is further phase shifted by a phase dot; and the resulting imaged intensity. The detected signal (histogram) has contributions from the operator expectation value (dashed line) as well as projection noise. (b) The first measurement creates excitations that are correlated with the measurement outcome and propagate for a time $\delta t$ prior to a second measurement.
@@
Generalized quantum measurements, however, allow for weak measurements that minimally disturb the system, but in exchange provide only limited information. These have long been associated with continuous monitoring of both open and closed quantum systems Guerlin et al. (2007); Minev et al. (2019), and are essential for closed loop quantum control Wiseman and Milburn (2009). Here, we make use of weak measurements in a many-body context, where they will allow the extraction of multi-time correlation functions, valid independent of quantum statistics and dimensionality. We specifically focus on homodyne weak measurement schemes [for example, realized by phase-contrast imaging (PCI) as shown in Fig. 1(a)] of atoms in a 1D lattice that report $n_{j}(t)$, a noisy estimate of number at lattice site $j$ density, and yield post-measurement wavefuntions that incorporate density fluctuations inferred from the measurement. The resulting excitations propagate for a time delay $\delta t$ prior to a second weak measurement of density giving $n_{j}(\delta t)$, as in Fig. 1(b). Intuitively the second measurement outcome includes a contribution from the system’s response to the first measurement, and we analytically confirm that the ensemble average  [^1] The ensemble can include trajectories for a single initial pure state, or different initial states sampled from a specified distribution. of such observations $\overline{n_{0}(0)n_{j}(\delta t)}$, converges to the real part of the un-equal time correlation function ${\rm Re}\left[\langle\hat{n}_{0}(0)\hat{n}_{j}(\delta t)\rangle\right]$. We numerically study an implementation of this approach for the 1D Bose-Hubbard (BH) model, and show that in the limit of vanishing measurement strength $S(q,\omega)$ obtained by the weak measurement process converges to that directly computed from the underlying wavefunctions. We conclude by showing that this method robustly captures long wavelength excitations even with realistic imaging limitations included.
@@
Homodyne detection– Our weak measurement protocol is based on homodyne detection, which could be experimentally implemented using PCI as schematically depicted in Fig. 1(a). When illuminated with coherent light of wavelength $\lambda$, an atomic ensemble in the object plane diffracts part of the probe laser by imprinting a position-dependent phase shift onto it. PCI is usually operated in the far-detuned limit, where absorption can be ignored, so the optical intensity just beyond the atomic ensemble is unchanged. By design, PCI is sensitive to changes in the phase quadrature: a phase dot at the Fourier plane of a Keplarian telescope phase shifts the unscattered light by a known phase value (typically $\pi/2$Altuntaş and Spielman (2021). In the image plane, the interference between scattered light and the phase-shifted probe changes the detected intensity by an amount proportional to the atomic density. The usual “photon shot noise” in this image is the projection noise associated with detecting the intensity of a coherent state. For a bright beam, this leads to Gaussian backaction on the atomic density.
@@
Analytical derivation—To be concrete, we consider bosonic atoms in a 1D lattice with native system Hamiltonian $\hat{H}_{\rm BH}$, describing the BH model, responsible for coherent dynamics. These dynamics are interjected by brief homodyne measurements of the atomic density, with durations short compared to the time scale for coherent evolution. Each measurement, at a discrete time $t$, gives an outcome
@@
in terms of the difference operator ${\delta\hat{n}_{j,t}}\equiv\hat{n}_{j}-\langle\hat{n}_{j}(t)\rangle$.
In our protocol [illustrated in Fig. 1(b)], we (1) perform an initial measurement at time $t=0$; (2) allow the system to undergo unitary evolution described by $\hat{U}(\delta t)=\exp\big(-i\hat{H}_{\rm BH}\delta t/\hbar\big{missing})$ for a time $\delta t$; and (3) perform a second measurement. We focus on the cross-correlation of the measurement results
@@
Terms such as $\overline{m_{j^{\prime},\delta t}\langle\hat{n}_{j,0}\rangle}$ are absent in this expression because the number at earlier times is uncorrelated with measurement noise at later times. By contrast, $\langle\hat{n}_{j^{\prime},\delta t}\rangle$ can be correlated with the noise $m_{j,0}$. We adopt a notation where expectation values with a prime
are with respect to the post measurement initial state. To lowest order, the expectation value of the number at time $\delta t$ is
$\langle\hat{n}_{j^{\prime}}(\delta t)\rangle^{\prime}$ $=\langle\hat{n}_{j^{\prime}}(\delta t)\rangle+\Gamma^{1/2}\sum_{j}m_{j,0}\bigg{[}\langle\hat{n}_{j}(0)\hat{n}_{j^{\prime}}(\delta t)\rangle$
$\ \ \ +\langle\hat{n}_{j^{\prime}}(\delta t)\hat{n}_{j}(0)\rangle-2\langle\hat{n}_{j}(0)\rangle\langle\hat{n}_{j^{\prime}}(\delta t)\rangle\bigg{]}.$
@@
as a function of spatial displacement $\delta j$, known as the Van Hove function Van Hove (1954) in x-ray and neutron scattering. This demonstrates our central observation: for $\Gamma\ll 1$, correlating subsequent measurement results yields time-separated correlation functions. This is a generic observation valid for pairwise combinations of weak measurements (potentially of different observables) with the general structure in Eqs. (1)-(2).
@@
We numerically obtain the ground state |$\ket{\psi(0)}$, and then initiate a specific trajectory by stochastically selecting a realization of the projection noise $m_{j,t=0}$. This allows us to compute the associated post-measurement state |$\ket{\psi^{\prime}(0)}$ using Eq. (2) and measurement outcome $n_{j,0}$. The next step is to evolve the resulting state according to $\hat{H}_{\rm BH}$ for a time $\delta t$, obtain $\langle\hat{n}_{j}(\delta t)\rangle$, select a second noise realization $m_{j,\delta t}$, and thereby obtain the second measurement outcome $n_{j,\delta t}$. We then average the correlation signal [Eq. (3)] over trajectories. To increase sampling efficiency in the numerical simulations, we use each initial noise realization $m_{j,t=0}$ to generate the full collection of second measurement outcomes, by computing $\langle\hat{n}_{j}(\delta t)\rangle$ and $m_{j,\delta t}$ at every desired measurement time. When reporting correlation functions, we further enhance the sampling efficiency by omitting the noise contribution to the second measurement, giving $n_{j,\delta t}=\langle\hat{n}_{j}(\delta t)\rangle$. The number of trajectories quoted in the figure captions reflects the number of distinct initial noise realizations.
Figure 2 plots the results of these simulations (see caption for simulation parameters); (a) and (b) show $\delta t$ and $\delta j$ cross-sections of the correlation function $G_{\delta j}(\delta t)$ with quantities extracted directly from MPS wavefunctions (black), and those obtained using Eq. (5) (red) agreeing within the statistical uncertainty (error bands). The duration of the MPS simulations are limited to the short times— $\delta t\times J\lesssim 3$ for current parameters—for which they are numerically exact (i.e., converged in truncation error per-timestep); this blocks numerical access to small $\omega$ when computing $S(q,\omega)$. Longer times (and therefore smaller $\omega$) are experimentally accessible, however, finite size effects will appear for large $\delta t$.
Panels (c) and (d) confirm that $S(q,\omega)$ extracted from the MPS wavefunction (left) is in correspondence with that obtained from the correlation signal (right). The key features of the excitation spectrum Fisher et al. (1989) are present: in the superfluid phase ($U/J=2$) a gapless phonon mode (linear for small $\omega$) is present; and in the Mott insulating phase ($U/J=4,5$) a energy gap clearly appears. In both cases the DSF contains sharp spectra features derived from well defined quasiparticle excitations: phonons in the superfluid phase giving way to particle-like excitations in the Mott phase.
@@
As briefly noted above, this approach can be be used to obtain more exotic two-time correlation functions of any pair of Hermitian operators that can be weakly measured. Such a process yields the expectation value of the ${\rm O}(\Gamma^{1/2})$ backaction operator of the first measurement [e.g., $\delta{\hat{n}}_{j}$ in Eq. (2)] multiplied by the measurement operator of the second measurement [e.g, $\hat{n}_{j^{\prime}}$ in Eq. (4)]. Examples include spin-spin correlation functions such as $\langle\hat{n}_{j,\uparrow}\hat{n}_{j^{\prime},\downarrow}\rangle$ in multi-component (or even multi-species) systems and even hybrid spatial-momentum correlations $\langle\hat{n}_{j}\hat{n}_{k}\rangle$ that combine the momentum space number density $\hat{n}_{k}$ with the real space density.

### v2

@@ Abstract
Much of our knowledge of quantum systems is encapsulated in the expectation value of Hermitian operators, experimentally obtained by averaging projective measurements. However, dynamical properties are often described by products of operators evaluated at different times; such observables cannot be measured by individual projective measurements, which occur at a single time. For example, the dynamical structure factor describes the propagation of density excitations, such as phonons, and is derived from the spatial density operator evaluated at different times. In equilibrium systems this can be obtained by first exciting the system at a specific wavevector and frequency, then measuring the response. Here, we describe an alternative approach using a pair of time-separated weak measurements, and analytically show that their cross-correlation function directly recovers the dynamical structure factor, for all systems, even far from equilibrium. This general schema can be applied to obtain the cross-correlation function of any pair of weakly observable quantities. We provide numerical confirmation of this technique with a matrix product states simulation of the one-dimensional Bose-Hubbard model, weakly measured by phase contrast imaging. We explore the limits of the method and demonstrate its applicability to real experiments with limited imaging resolution.
@@
The properties of many body quantum systems are often encoded in response functions that quantify the reaction of the system to weak external perturbations. The dynamical structure factor (DSF) $S(q,\omega)$, describing a system’s response to density perturbations with wavenumber $q$ and frequency $\omega$, derives from the expectation value $\langle\hat{n}(0,0)\hat{n}(x,t)\rangle$ of the spatial density operator $\hat{n}(x,t)$ at different times and positions Ashcroft and Mermin (1976). For equilibrium systems, such quantities are experimentally obtained by observing the system’s response to a suitable perturbation. For example neutron scattering gives access to $S(q,\omega)$ (possibly in a spin-dependent way) in materials Sturm (1993), where it has contributed to the understanding of high temperature superconductivity Hinkov et al. (2007), and topological spin systems Yao et al. (2018); Banerjee et al. (2018) to name a few. Analogously for ultracold atoms, $S(q,\omega)$ can be obtained by Bragg scattering far-detuned laser light off atomic ensembles Birkl et al. (1995); this has shed light on weak- and strongly-interacting Bose-Einstein condensates Steinhauer et al. (2002); Pino et al. (2011), structural phase transitions Landig et al. (2015), and unitary Fermi gases Biss et al. (2022). The spectral function, a related response function correlating fields rather than densities, can be measured using similar techniques Brown et al. (2020); Bohrdt et al. (2020, 2021). By contrast, we focus on measuring density-density correlations $\langle\hat{n}(0,0)\hat{n}(x,t)\rangle$, and thereby $S(q,\omega)$, by weak quantum measurements alone Xu et al. (2015).
Unlike the methods described above, this technique is not limited to equilibrium systems, and is applicable to any state, pure or mixed, undergoing Hamiltonian dynamics. This connects to previous work Hung et al. (2011), that showed the static structure factor $S(q)$ can be obtained from the same-time density-density correlation function $\langle\hat{n}(0)\hat{n}(x)\rangle$ computed from simple projective measurements of density. Because projective measurements collapse the wavefunction at a well-defined time, they prevent access to the two-time correlations we require.
@@
Generalized quantum measurements, however, allow for weak measurements that minimally disturb the system, but in exchange provide only limited information. These have long been associated with continuous monitoring of both open and closed quantum systems Guerlin et al. (2007); Minev et al. (2019), and are essential for closed loop quantum control Wiseman and Milburn (2009). Here, we make use of weak measurements in a many-body context, where they will allow the extraction of multi-time correlation functions, valid independent of quantum statistics and dimensionality. We specifically focus on homodyne weak measurement schemes [for example, realized by phase-contrast imaging (PCI) as shown in Fig. 1(a)] of atoms in a 1D lattice that report $n_{j}(t)$, a noisy estimate of number at lattice site $j$ density, and yield post-measurement wavefunctions that incorporate density fluctuations inferred from the measurement. The excitations resulting from the first weak measurement at time $t=0$ propagate for a time delay $\delta t$ prior to a second weak measurement of density giving $n_{j^{\prime}}(\delta t)$, as in Fig. 1(b). Intuitively the second measurement outcome includes a contribution from the system’s response to the first measurement, and we analytically confirm that the ensemble averaged  [^1] The ensemble can include trajectories for a single initial pure state, or different initial states sampled from a specified distribution. cross-correlations between these observations $\overline{n_{j}(0)n_{j^{\prime}}(\delta t)}$, converges to the real part of the un-equal time expectation value: ${\rm Re}\left[\langle\hat{n}_{j}(0)\hat{n}_{j^{\prime}}(\delta t)\rangle\right]$. This interpretation is similar to that presented Ref. Xu et al., 2015 in the context of photodetection theory.
We numerically study an implementation of this approach for the 1D Bose-Hubbard (BH) model, and show that in the limit of vanishing measurement strength $S(q,\omega)$ obtained by the weak measurement process converges to that directly computed from the underlying wavefunctions. We conclude by showing that this method robustly captures long wavelength excitations even with realistic imaging limitations included.
@@
Homodyne detection– Our weak measurement protocol is based on homodyne detection, which could be experimentally implemented using PCI as schematically depicted in Fig. 1(a). When illuminated with coherent light of wavelength $\lambda$, an atomic ensemble in the object plane diffracts part of the probe laser by imprinting a position-dependent phase shift onto it. PCI is usually operated in the far-detuned limit, where absorption can be ignored, so the intensity of the probe laser just after having traversed the atomic ensemble is unchanged. By design, PCI is sensitive to changes in the phase quadrature: a phase dot at the Fourier plane of a Keplerian telescope phase shifts the unscattered light by a known phase value (typically $\pi/2$) while leaving the scattered light unchanged Altuntaş and Spielman (2021). PCI is an interferometric method in which the unscattered probe serves as the reference beam (i.e., local oscillator), while the scattered component carries information about the atomic ensemble. In the image plane, the interference between scattered light and the phase-shifted probe changes the detected intensity by an amount proportional to the atomic density. The usual “photon shot noise” in this image is the projection noise associated with detecting the intensity of a coherent state. For a bright beam, this leads to backaction described by a Kraus operator that is a Gaussian function of the atomic density operator.
Figure 1: System and measurement schema. (a) PCI as an example of homodyne detection. Probe light is first phase shifted by the atomic ensemble (light scattered by more than the acceptance angle $\theta$ is not detected); the un-scattered light is further phase shifted by a phase dot; and the resulting interference between the scattered and the un-scattered light is imaged. The detected signal (histogram) has contributions from the operator expectation value (dashed line) as well as projection noise. (b) A pair of time-separated weak measurements. The first measurement creates excitations that are correlated with the measurement outcome and propagate for a time $\delta t$ prior to a second measurement.
@@
Analytical derivation—To be concrete, we consider bosonic atoms in a 1D lattice with an arbitrary native system Hamiltonian $\hat{H}$, responsible for coherent dynamics. These dynamics are interjected by brief homodyne measurements of the atomic density, with durations short compared to the time scale for coherent evolution. Each measurement, at a discrete time $t$, gives an outcome
@@
in terms of the difference operator ${\delta\hat{n}_{j,t}}\equiv\hat{n}_{j}-\langle\hat{n}_{j}(t)\rangle$. In following expressions we adopt a notation where expectation values with a prime
are with respect to the post measurement initial state.
In our protocol [illustrated in Fig. 1(b)], we (1) perform an initial measurement at time $t=0$; (2) allow the system to undergo unitary evolution described by $\hat{U}(\delta t)=\exp\big(-i\hat{H}\delta t/\hbar\big{missing})$ for a time $\delta t$; and (3) perform a second measurement. We focus on the cross-correlation of the measurement results
@@
Terms such as $\overline{m_{j^{\prime},\delta t}\langle\hat{n}_{j}(0)\rangle}$ are absent in this expression because the number at earlier times is uncorrelated with measurement noise at later times. By contrast, $\langle\hat{n}_{j^{\prime}}(\delta t)\rangle$ can be correlated with the noise $m_{j,0}$. Both terms in Eq. (4) have clear physical meaning: the first term identifies excitations created by measurement backaction, while the second term is simply the correlation between the initial density and the averaged final density that one might naively expect.
To lowest order in $\Gamma$, the expectation value of the number at time $\delta t$ is
$\langle\hat{n}_{j^{\prime}}(\delta t)\rangle^{\prime}$ $=\langle\hat{n}_{j^{\prime}}(\delta t)\rangle+\Gamma^{1/2}\sum_{j}m_{j,0}\big{[}\langle\hat{n}_{j}(0)\hat{n}_{j^{\prime}}(\delta t)\rangle$
$\ \ \ +\langle\hat{n}_{j^{\prime}}(\delta t)\hat{n}_{j}(0)\rangle-2\langle\hat{n}_{j}(0)\rangle\langle\hat{n}_{j^{\prime}}(\delta t)\rangle\big{]}.$ (5)
@@
as a function of spatial displacement $\delta j$, known as the Van Hove function Van Hove (1954) in x-ray and neutron scattering. This demonstrates our central observation: for $\Gamma\ll 1$, correlating subsequent measurement results yields time-separated correlation functions (the appendix comments on the consequence of stronger measurements.). This is a generic observation valid for pairwise combinations of weak measurements (potentially of different observables) with the general structure in Eqs. (1)-(2).
@@
We numerically obtain the ground state |$\ket{\psi(0)}$, and then initiate a specific trajectory by stochastically selecting a realization of the projection noise $m_{j,t=0}$. This allows us to compute the associated post-measurement state |$\ket{\psi^{\prime}(0)}$ using Eq. (2) and measurement outcome $n_{j,0}$. The next step is to evolve the resulting state according to $\hat{H}_{\rm BH}$ for a time $\delta t$, obtain $\langle\hat{n}_{j}(\delta t)\rangle$, select a second noise realization $m_{j,\delta t}$, and thereby obtain the second measurement outcome $n_{j,\delta t}$. We then average the correlation signal [Eq. (4)] over trajectories. To increase sampling efficiency in the numerical simulations, we use each initial noise realization $m_{j,t=0}$ to generate the full collection of second measurement outcomes, by computing $\langle\hat{n}_{j}(\delta t)\rangle$ and $m_{j,\delta t}$ at every desired time increment. When reporting correlation functions, we further enhance the sampling efficiency by omitting the noise contribution to the second measurement, giving $n_{j,\delta t}=\langle\hat{n}_{j}(\delta t)\rangle$. The number of trajectories quoted in the figure captions reflects the number of distinct initial noise realizations.
Figure 2 plots the results of these simulations (see caption for simulation parameters); (a) and (b) show $\delta t$ and $\delta j$ cross-sections of the correlation function $G_{\delta j}(\delta t)$ with quantities extracted directly from MPS wavefunctions (black), and those obtained using Eq. (6) (red) agreeing within the statistical uncertainty (error bands). The duration of the MPS simulations are limited to the short times— $\delta t\times J\lesssim 3$ for current parameters—for which they are numerically exact (i.e., converged in truncation error per-timestep); this blocks numerical access to small $\omega$ when computing $S(q,\omega)$. Longer times (and therefore smaller $\omega$) are experimentally accessible, however, finite size effects will appear for large $\delta t$. For example, at $U/J=2$ the correlation signal nominally spreads by approximately two lattice sites per time increment of $\delta t\times J$. Then, in a 51-site 1D BH chain, an excitation initially located at the center of the system will reach the boundary in $\approx 12$ tunneling times.
Panels (c) and (d) confirm that $S(q,\omega)$ extracted from the MPS wavefunction (left) is in correspondence with that obtained from the correlation signal (right). The key features of the excitation spectrum Fisher et al. (1989) are present: in the superfluid phase ($U/J=2$) a gapless phonon mode (linear for small $\omega$) is present; and in the Mott insulating phase ($U/J=4,5$) a energy gap clearly appears. In both cases the DSF contains sharp spectra features derived from well defined quasiparticle excitations: phonons in the superfluid phase giving way to particle-like excitations in the Mott phase. The artifacts present in the correlations signal in Fig. 2(d) arise from a hybrid term that combines both statistical and systematic uncertainties (see Appendix A.).
@@
As briefly noted above, this approach can be be used to obtain two-time correlation functions of any pair of Hermitian operators that can be weakly measured. Such a process yields the expectation value of the $\mathrm{O}(\Gamma^{1/2})$ backaction operator of the first measurement [e.g., $\delta{\hat{n}}_{j}$ in Eq. (2)] multiplied by the measurement operator of the second measurement [e.g, $\hat{n}_{j^{\prime}}$ in Eq. (3)]. As a straightforward example, in spinor systems quantities such the “spin Van Hove function” $\langle(\hat{n}_{\uparrow,j,0}-\hat{n}_{\downarrow,j,0})(\hat{n}_{\uparrow,j^{\prime},\delta t}-\hat{n}_{\downarrow,j^{\prime},\delta t})\rangle$ (essentially correlations in the $\hat{\sigma}_{z}$ magnetization) could be easily measured by performing PCI with the probe laser red-detuned with respect to one spin state and blue-detuned with respect to the other Shin et al. (2006). Measuring $\sigma_{x}$ and $\sigma_{y}$ magnetization requires additional coherent control to transform the desired quantities to and from the measurement basis. Other examples include spin-spin cross-correlation functions such as $\langle\hat{n}_{j,\uparrow}\hat{n}_{j^{\prime},\downarrow}\rangle$ in multi-component (or even multi-species) systems and even hybrid spatial-momentum correlations $\langle\hat{n}_{j}\hat{n}_{k}\rangle$ that combine the momentum space number density $\hat{n}_{k}$ with the real space density.